# Scalable Oversight for Superhuman AI via Recursive Self-Critiquing

### v2

@@ 5.2 Experimental Results
We plot the average imporvement of critique and recursive cirtics over response stage in Figure 2, where we find that the performance of critiques and higher-order critiques remains constrained in most scenarios. In detail, critique, critique of critique ($C^{2}$) and third-order critiques ($C^{3}$) struggle to surpass the accuracy levels achieved at the Response stages. This phenomenon suggests that current models may have limited capacity for self-critiquing, which is also partially validated in the work of Huang et al. (2023),  Tang et al. (2025) and summarized by Kamoi et al. (2024). This finding further highlight the importance of post-training models to perform better critique tasks as in McAleese et al. (2024).
@@
Input Type Gemma2 9B Instruct Qwen2.5 14B Instruct
@@
We further compare the performance of Qwen and Gemma model across different datasets, following the setting in Section 4.2. The results are shown in Table 4 and Table 5. From these results, we can observe that the disparity of higher-order critiquing ability exists in specific models. Although performance is limited, Qwen-2.5-14B-Instruct exhibits more effective in high-order critiquing, as we can observe progressive improvements between response and recursive critiques across stages. Further performance gap analysis reveals that the main discrepancies may arise from those gap in distinguishing the true statements from mixed true and false inputs (Table 6).
Figure 3: Illustration of challenges in RLHF: (1) Golden RM has fixed capability ceiling; (2) Proxy RM suffers from degrading supervision as policies improve; (3) Iterative annotation provides temporary improvements through human intervention; (4) Tool-augmentation achieves higher but still limited performance. (5) The ideal scenario requires RM capabilities to scale alongside policy model improvements.
6 Related Works
Many works have been dedicated to developing protocols that enable weak annotators to supervise strong AI systems.
Debate
This protocol (Irving et al., 2018a) involves two models/humans debating opposing answers. Arguments can be generated synchronously or asynchronously. In certain variants like iteractivate debate, the Judge can interact with the Debaters (Khan et al., 2024). Several studies have analyzed debate as a scalable oversight approach in practical settings. For example, Khan et al. (2024) and Michael et al. (2023) conducted machine and human experiments respectively on QuALITY dataset with promising results, while Kenton et al. (2024a) evaluated debate extensively on broader reasoning tasks but found limited benefits. The key difference between critique of critique and Debate is that critique does not assume a zero-sum game - the Judge can reject both sides and make independent judgments, or accept consensus answers from both response or critiques.
Task Decomposition
This approach posits that oversight of difficult problems can be reduced to supervising and combining solutions to sub-problems. Previous works  (Christiano et al., 2018; Wu et al., 2021) have explored this concept for both policy models and reward models (i.e., Recursive Reward Modeling). While critique of critique can be viewed as reducing complex tasks to specific sub-tasks, it does not require solving all sub-problems to resolve the overall task. Its recursive structure resembles debate more than task decomposition - it employs depth-first rather than breadth-first search in breaking down problems.
7 Discusssion
Limitations in Current AI Alignment Methods.
RLHF has emerged as the dominant approach in AI alignment, building upon the fundamental principle that “verification is easier than generation” (Irving et al., 2018b). As illustrated in Figure 3, the optimal RLHF setup requires direct human preferences (Golden RM) for policy optimization. However, the acquisition of real-time human feedback during training presents significant operational challenges. This limitation necessitates the deployment of static reward models as proxies for human judgment. Such reliance on static proxies introduces a fundamental challenge: reward hacking (Gao et al., 2022; Karwowski et al., 2023), optimizing against proxy reward models rather than ideal human preferences can lead to policies that diverge from intended objectives due to Goodhart’s Law (Manheim & Garrabrant, 2019; Karwowski et al., 2023; Wen et al., 2024). While approaches such as iterative annotation and tool augmentation (Li et al., 2024a; Gou et al., 2024) provide intermediate solutions, they ultimately encounter the supervision capability limitations. Addressing reward hacking requires reward model capabilities to scale alongside policy model improvements. The recursive criticism framework, while not eliminating reward hacking entirely, offers a promising approach by enabling sustained human oversight even as direct evaluation becomes intractable.
The Possible Mechanism of Recursive Self-Critiquing.
The effectiveness of recursive self-critiquing in Human-Human and Human-AI experiments likely stems from several key mechanisms. First, higher-order criticism progressively shifts human’s attanetion from specific details to more abstract evaluation principles, making the task more tractable even as context length increases. Second, each critique level provides structured context that helps frame subsequent evaluation, allowing critiques to build on previous analyses rather than starting fresh. Third, the recursive structure transforms absolute evaluation tasks into a series of pairwise judgments, leveraging humans’ cognitive advantage in relative assessment over absolute evaluation. This aligns with cognitive science findings (Jones & Inglis, 2015; Kelly et al., 2022) that humans often perform better at comparative judgment than absolute evaluation, particularly when provided with structured information.
Improving AI Critique Capabilities.
From AI-AI experiments, our findings suggest several promising directions for training AI systems with recursive critique capabilities. Firstly, models could be explicitly trained to more accurately identify critical errors (Xi et al., 2024), rather than merely aggregating lower-level judgments. Additionally, future work can explore ways to maintain sufficient diversity in AI-generated critiques, perhaps through ensemble methods or explicit diversity-oriented training objectives, to better mirror the effectiveness seen in human recursive self-critiquing experiments. Finally, research into the optimal number of recursive levels and the diminishing returns of additional critique iterations could help establish practical guidelines for implementing recursive critique systems effectively.
8 Conclusion
This work investigates how to obtain reliable supervision signals when AI capabilities surpass human abilities. Through comprehensive experiments in human-human, human-AI, and AI-AI contexts, we examine the hypotheses that critique of critique is easier than critique and demonstrate that this difficulty relation holds recursively. These insights of recursive self-critiquing mechanisms could be crucial for maintaining effective oversight in scenarios where direct human evaluation becomes infeasible, and suggest a promising pathway for scalable oversight.
Impact Statement
This recursive self-critiquing framework aims to address challenges in scalable AI oversight, enabling more effective human supervision as AI systems become increasingly capable. Beyond technical contributions, this research promotes responsible AI development by involving more people from different backgrounds in AI supervision. While acknowledging potential limitations, we believe this work helps promote social welfare via contributing scalable oversight mechanisms.
Appendix A Human Experiments Guidelines
A.1 Response Stage
Recommmanded Annotation Template
  • Option B is correct because [evidence + reasoning].
{CJK*}
UTF8gkai
Table 7: High quality and low quality response examples.
Quality Definition Type Example English Translation
High quality Contains three elements: textual evidence, reasoning, and conclusion. Clear and coherent expression with logical flow. English 根据题中的”before the end of the century”可定位到原文”Scientists have already pointed out that unless something … before this century is out”。从中可以得知如果不采取措施限制人口快速增长或开发新的食物来源,数百万人将在本世纪结束前死于饥饿。因此可推断作者认为世界最大的问题是如何养活日益增长的人口,选B。 Based on the phrase ”before the end of the century”, we can locate ”Scientists have already pointed out that unless something … before this century is out”. This indicates that without measures to limit population growth or develop new food sources, millions will face starvation. Therefore, feeding the growing population appears to be the major challenge, supporting option B.
Chinese 文章第三段说:”由于杂交水稻不同熟期组合的出现,全国各地涌现出各种与杂交水稻种植相配套的新型种植模式。”杂交水稻和新型种植模式的出现是因果关系,而不是正好与新型种植模式相配,所以选D。 The third paragraph states: ”Due to the emergence of hybrid rice varieties with different maturity periods, new planting patterns have emerged nationwide to match hybrid rice cultivation.” The relationship between hybrid rice and new planting patterns is causal, not just coincidental matching, therefore D is correct.
Math 首先化简$f(x)=2\cos^{2}x-\sin^{2}x+2$,根据二倍角公式$\cos 2x=2\cos^{2}x-1$,得到$12\cos^{2}x=\cos 2x+$ $\sin^{2}x+\cos^{2}x=1$,所以$\sin^{2}x=(1-\cos 2x)/2$。最终得到$f(x)=\frac{3}{2}\cos 2x+\frac{5}{2}$。通过周期计算和最值分析,得到答案B。 First simplify $f(x)=2\cos^{2}x-\sin^{2}x+2$. Using double angle formula $\cos 2x=2\cos^{2}x-1$, we get $12\cos^{2}x=\cos 2x+$ 1. Since $\sin^{2}x+\cos^{2}x=1$, we derive $\sin^{2}x=(1-\cos 2x)/2$. Finally $f(x)=\frac{3}{2}\cos 2x+\frac{5}{2}$. Through period calculation and maximum analysis, we arrive at answer B.
Low quality Missing key elements, unclear reasoning, or lack of evidence support. English 文章第一句”The gift of being able to describe a face accurately is a rare one”就点明文章主要内容为A。 The first sentence ”The gift of being able to describe a face accurately is a rare one” directly points to option A.
Chinese 答案C错在:那些已经被认定,应…,原文说的是这种代代相传的非物质文化遗产得到创新(过程中),同时使他们自己具有一种认同感和历史感。 Option C is wrong because: those already recognized should…, the text actually discusses how this inherited intangible cultural heritage achieves innovation while maintaining a sense of identity and history.
[…]
Quality Requirements
Response annotations must satisfy four fundamental criteria:
  • Relevance: Direct connection to the question and source text
  • Organization: Clear logical structure and information flow
  • Clarity: Concise expression without unnecessary complexity
  • Coherence: Smooth transitions between reasoning steps
A.2 Critique Stage Annotation
Recommended Annotation Template
{CJK*}
UTF8gkai
Table 8: High quality and low quality critique annotation examples.
Question Response1 Response2
High Quality Examples
认可Response 2的答案(B),他引用的文章内容「饮食起居所需要的陶罐、簸箕、箩筐」和选项B「医药的陶罐、竹器等物品」确实是不一致的。而Response1的选择(A),他引用[农业耕作所需要的犁、耙、锄,饮食起居所需要的陶罐、簸箕、箩筐,…都留下了手工艺的印记]里的”手工艺”我认为是传统技艺的一部分,因此不应该选答案A。因此答案应该是(B)。
Low Quality Examples
认可response 1的答案,不同意Response2的答案,答案应该是(B)。
Question Response1 Response2
According to the text, which of the following does NOT reflect the inheritance value of traditional craftsmanship? A. The use of agricultural tools like plows, harrows, and hoes B. The making of medical pottery and bamboo utensils in daily life C. The handicrafts in social industries such as papermaking, printing, and dyeing D. The folk art creations such as paper-cutting, shadow puppets, and New Year paintings The author mainly introduces traditional crafts as intangible cultural heritage with rich cultural connotations and artistic value. Traditional crafts possess livelihood value, artistic value, and historical-cultural value as their inheritance value. Option A does not reflect traditional craft inheritance value. While the text mentions the tools like plows, harrows, and hoes as traditional handicraft marks, reflecting craft inheritance value, the option focuses on tool usage, which is not demonstrated. The relevant content states [Agricultural tools like plows, harrows, and hoes, kitchenware like pottery, baskets, and hampers, transportation tools like carriers, boats, and sedan chairs, and daily necessities like brocade, dyed cloth, and paper, all bear handicraft marks]. Therefore, I believe the answer is ”A”. The article mainly discusses the definition, value, and importance of traditional crafts. For Option A: The text mentions ”agricultural tools like plows, harrows, and hoes” and their ”livelihood value”, suggesting these tools reflect traditional craft inheritance value. Additionally, Option C is supported by ”daily necessities like brocade, dyed cloth, and paper” showing ”livelihood and artistic value”, while Option D is evidenced by ”wood carving, stone carving, jade carving, clay sculpture, and dough modeling bring aesthetic pleasure, while paper-cutting, paper carving, shadow puppets, and New Year paintings have rich cultural connotations and artistic value”. Option B is incorrect as the text mentions ”kitchenware pottery” rather than medical pottery. Based on this analysis, the answer is ”B”.
High Quality Examples
Disagree with Response 1’s answer (A). While Response 1 cites relevant passages, its interpretation that ”focus is on tool usage” misses the point about traditional craft inheritance value. The cited phrase ”left craftmanship marks” actually demonstrates that agricultural tools reflect traditional craft value. Agree with Response 2’s answer (B). Response 2 correctly analyzes Option B, noting that the text mentions ”kitchenware pottery” rather than ”medical pottery”, with accurate textual evidence. They also provide well-supported analysis for Options A, C, and D. However, their reasoning for Option A needs correction - the connection between ”agricultural tools” and ”livelihood value” isn’t supported by being in different paragraphs. Nevertheless, the fourth paragraph’s opening about ”handicrafts beginning with the use of natural materials and tool manufacturing” supports that Option A reflects traditional craft value. Therefore, the answer should be (B).
Agree with Response 2’s answer (B). Their citation of ”kitchenware pottery” from the text clearly contradicts Option B’s ”medical pottery”. Regarding Response 1’s choice of (A), the reference to ”handicraft marks” in the passage about agricultural tools suggests this is part of traditional craftsmanship, so A should not be the answer. Therefore, the answer is (B).
Support Response 2’s answer (B). They correctly identified the discrepancy between ”kitchenware pottery” in the text and ”medical pottery” in Option B, which appears to be an intentional distinction. While Response 1 analyzes why Option A might be incorrect, based on the analysis of Option B being clearly wrong, Option A cannot be the answer. Therefore, the answer should be (B).
Low Quality Examples
Quality Requirements
critique annotations must satisfy five fundamental criteria:
  • Relevance: Direct connection to the question and source text
  • Organization: Clear logical structure and information flow
  • Clarity: Concise expression without unnecessary complexity
  • Coherence: Smooth transitions between reasoning steps
  • Objectivity: Fair analysis of responses’ strengths and weaknesses
A.3 Higher-Order Critique Stage
Recommended Annotation Template
Quality Requirements
  • Relevance: Direct connection to the question and critics’ analyses.
  • Organization: Clear logical structure and information flow.
  • Clarity: Concise expression without unnecessary complexity.
  • Coherence: Smooth transitions between reasoning steps.
  • Objectivity: Fair analysis of critics’ strengths and weaknesses.
Examples of high-quality and low-quality higher-order critiques are presented in Tables 10 and 11.
{CJK*}
UTF8gkai
Question Response 1 Response 2
Critic 1 Critic 2
不同意Response1的答案(A)。Response1引用了相关的段落,但认为相关内容「重点在于对工具的利用」,而不是体现传统技艺传承价值。但他引用的内容里提到了「留下了手工艺的印记」,我认为这个内容正体现了农业生产的犁、耙、锄是体现传统技艺传承价值的。 Response2分析了选项B,并指出「原文提到的是饮食起居的陶罐,而不是医药的陶罐」,他引用了文章中对应内容确实和选项是不符合的。此外,他还分析了选项ACD,并找到了对应的参考内容佐证。因此答案应该是(B)。 认可Response 2的答案(B),他引用的文章内容「饮食起居所需要的陶罐、簸箕、箩筐」和选项B「医药的陶罐、竹器等物品」确实是不一致的。而Response1的选择(A),他引用[农业耕作所需要的犁、耙、锄,饮食起居所需要的陶罐、簸箕、箩筐,…都留下了手工艺的印记]里的”手工艺”我认为是传统技艺的一部分,因此不应该选答案A。因此答案应该是(B)。
High Quality Examples
Low Quality Examples
Critc 1/2的答案是对,应该是(B)。
Question Response 1 Response 2
According to the text, which of the following does NOT reflect the inheritance value of traditional craftsmanship? A. The use of agricultural tools like plows, harrows, and hoes B. The making of medical pottery and bamboo utensils in daily life C. The handicrafts in social industries such as papermaking, printing, and dyeing D. The folk art creations such as paper-cutting, shadow puppets, and New Year paintings The author mainly introduces traditional crafts as intangible cultural heritage with rich cultural connotations and artistic value. Traditional crafts possess livelihood value, artistic value, and historical-cultural value as their inheritance value. Option A does not reflect traditional craft inheritance value. While the text mentions tools like plows, harrows, and hoes as traditional handicraft marks, reflecting craft inheritance value, the option focuses on tool usage, which is not demonstrated. The relevant content states [Agricultural tools like plows, harrows, and hoes, kitchenware like pottery, baskets, and hampers, transportation tools like carriers, boats, and sedan chairs, and daily necessities like brocade, dyed cloth, and paper, all bear handicraft marks]. Therefore, I believe the answer is ”A”. The article mainly discusses the definition, value, and importance of traditional crafts. For Option A: The text mentions ”agricultural tools like plows, harrows, and hoes” and their ”livelihood value”, suggesting these tools reflect traditional craft inheritance value. Additionally, Option C is supported by ”daily necessities like brocade, dyed cloth, and paper” showing ”livelihood and artistic value”, while Option D is evidenced by ”wood carving, stone carving, jade carving, clay sculpture, and dough modeling bring aesthetic pleasure, while paper-cutting, paper carving, shadow puppets, and New Year paintings have rich cultural connotations and artistic value”. Option B is incorrect as the text mentions ”kitchenware pottery” rather than medical pottery. Based on this analysis, the answer is ”B”.
Critic 1 Critic 2
Disagree with Response 1’s answer (A). While Response 1 cites relevant passages, its interpretation that ”focus is on tool usage” misses the point about traditional craft inheritance value. The cited phrase ”left craftmanship marks” actually demonstrates that agricultural tools reflect traditional craft value. Response 2 correctly analyzes Option B, noting that the text mentions ”kitchenware pottery” rather than ”medical pottery”, with accurate textual evidence. They also provide well-supported analysis for Options A, C, and D. Therefore, the answer should be (B). Agree with Response 2’s answer (B). Their citation of ”kitchenware pottery” from the text clearly contradicts Option B’s ”medical pottery”. Regarding Response 1’s choice of (A), the reference to ”handicraft marks” in the passage about agricultural tools suggests this is part of traditional craftsmanship, so A should not be the answer. Therefore, the answer is (B).
High Quality Examples
Agree with both Critics’ answer (B). Both critics point out that the discrepancy between ”kitchenware pottery” in the text and ”medical pottery” in Option B shows it does not reflect traditional craft inheritance value.
Agree with both Critics’ analysis of option B, noting the clear difference between ”kitchenware pottery” in the text and ”medical pottery” in the option. However, Critic 2’s reasoning about Response 1’s option A analysis is insufficient - ”handicraft marks” doesn’t necessarily equate to ”traditional crafts”, though this doesn’t affect the final judgment as option B is clearly correct.
Low Quality Examples
Critics 1/2 are correct, the answer should be (B).
Appendix B AI-AI Experiment Details
This section introduces the specific setup of our AI-AI experiments. We employ a unified sampling strategy to evaluate multiple models across different tasks. Detailed explanations are provided in the following sections.
B.1 Tasks Info
Our experiments are conducted across five types of tasks. Below is a detailed introduction to each task category.
  • MATH(Hendrycks et al., 2021b) is a mathematical problem-solving dataset consisting of 12,500 challenging competition-level math problems, designed to assess machine learning models’ mathematical reasoning abilities. Each problem is accompanied by a fully worked-out step-by-step solution, enabling models to learn how to generate answer derivations and explanations.
  • GPQA(Rein et al., 2023) is a highly challenging multiple-choice question dataset consisting of 448 questions crafted by domain experts in biology, physics, and chemistry. The dataset is designed to assess the reasoning capabilities of both human experts and state-of-the-art AI models on complex scientific topics. To ensure its difficulty and quality, questions were validated by experts with PhD-level knowledge, achieving an accuracy of only 65% (or 74% after correcting clear retrospective mistakes). In contrast, highly skilled non-expert validators, even with unrestricted web access for over 30 minutes per question, achieved only 34% accuracy.
  • TruthfulQA(Lin et al., 2022b) evaluates the truthfulness of language models in answering questions, comprising 817 questions across 38 categories, including health, law, finance, and politics. The questions were carefully designed to reflect common human misconceptions or false beliefs, making them particularly challenging. To perform well, models must avoid generating false answers learned from imitating human-written text, which often contains misinformation.
  • BoolQ(Clark et al., 2019) is a reading comprehension dataset designed to study naturally occurring yes/no questions, meaning questions that arise spontaneously in unprompted and unconstrained settings. The dataset presents unexpected challenges, as its questions often involve complex, non-factoid information and require entailment-like inference rather than simple fact retrieval.
  • MMLU-Pro(Wang et al., 2024a) is an enhanced version of MMLU designed to go beyond MMLU’s primarily knowledge-driven evaluation. MMLU-Pro incorporates more challenging reasoning-focused questions, expands the answer choice set from 4 to 10 options, and removes trivial and noisy questions from MMLU. Experimental results show that MMLU-Pro significantly increases difficulty, leading to an accuracy drop of 16% to 33% compared to MMLU.
B.2 Sampling Strategy
In the AI-to-AI experiment, to ensure fairness across different stages of effort, we followed the sampling strategy illustrated in Figure 4 to sample the responses of the model to a question and critics of various orders. Each sampling begins by obtaining 7 responses to the same question. From the first two responses, we further derive 5 critics. Similarly, we generate 3 critics of critics and 1 critics of critics of critics. To ensure the reliability of the results, we repeat the entire process 10 times for the same question and report the average outcomes of these ten iterations. To enhance the diversity of the sampling process, we set the sampling temperature to 1 and top-p to 0.95. The prompts used at different sampling stages are listed in the Appendix B.4.
Figure 4: The Sampling Strategy of AI Self Recursive Critiquing.
B.3 Additional Results
In this section, we provide additional results from AI-AI experiments. Tables 13 and 13 include the results from Gemma2 9B Instruct and Qwen-2.5 14B Instruct on MMLU-Pro and BoolQ.
@@ B.3 Additional Results
B.4 Generation Prompts
In this study, we explore whether AI can engage in Self Recursive Critiquing in an inference-only scenario. We employ the structured prompt illustrated in Figure 5, 6, 7, 8, 9 to obtain consistent forms of response, critique, and higher-order critique across different models and datasets. Given the variations in how different models adhere to and comprehend instructions, the prompt structure is slightly adjusted for each model. These adjustments primarily focus on constraints related to output length and the format of the decision-making answers.
@@ B.4 Generation Prompts
Prompt for $C^{2}$ Generation You are given a multiple-choice question. And two individuals, Person 1 and Person 2, have selected their answers and provided their explanations for their choices. Additionally, two more individuals, Reviewer 1 and Reviewer 2, have read these explanations and provided their evaluations of Person 1’s and Person 2’s reasoning. Your task is to identify which answer is correct based on their explanations and the information known about the question. Follow this structure for your response: - Explanation of Choice: Compare both explanations to your knowledge about the topic and determine which aligns better with the correct answer, starts with ‘Explanation: ‘. - Analysis of Other Options: Review the explanation provided by each person. Evaluate the reasoning and evidence behind each choice and point out any inaccuracies or correct assumptions. - Answer: On a separate line, starts with ‘Answer: ‘, state your chosen option (A, B, C, or D) only, without any additional text. ### Question: {question} ### Options: {options} ### Person 1’s Response: {gen1} ### Person 2’s Response: {gen2} ### Reviewer 1’s Response: {c1} ### Reviewer 2’s Response: {c2} Example Input: ### Question: Which element has the atomic number 6? ### Options: A) Nitrogen  B) Oxygen  C) Carbon  D) Helium ### Person 1’s Response: Chosen Answer: C Explanation: Carbon is the element with atomic number 6, well-known for being the basis of organic chemistry. ### Person 2’s Response: Chosen Answer: A Explanation: Nitrogen is important for life on Earth, making up a large portion of the atmosphere. ### Reviewer 1’s Response: Explanation: Person 1 accurately states that Carbon has the atomic number 6. This explanation is correct and directly addresses the question. Carbon is indeed a fundamental element … Answer: C ### Reviewer 2’s Response: Explanation: Person 1 has correctly identified Carbon as the element with atomic number 6. Their … Answer: C Example Output: Explanation: - Reviewer 1 provides a clear and direct evaluation of Person 1’s and Person 2’s responses. They … Answer: C
Figure 7: AI generartion template in $C^{2}$ Stage

### v3

@@ C.2 Experimental Results
We plot the average improvement of critique and recursive cirtics over response stage in Figure 9, where we find that the performance of critiques and higher-order critiques remains constrained in most scenarios. In detail, critique, critique of critique ($C^{2}$) and third-order critiques ($C^{3}$) struggle to surpass the accuracy levels achieved at the Response stages. This phenomenon suggests that current models may have limited capacity for self-critiquing, which is also partially validated in the work of Huang et al. [2023],  Tang et al. [2025] and summarized by Kamoi et al. [2024]. This finding further highlight the importance of post-training models to perform better critique tasks as in McAleese et al. [2024]. Tables 13 and 13 include the results from Gemma2 9B Instruct and Qwen-2.5 14B Instruct on MMLU-Pro and BoolQ.
@@
Input Type Gemma2 9B Qwen2.5 14B
@@
We further compare the performance of Qwen and Gemma model across different datasets, following the setting in Section 4.2. The results are shown in Table 10 and Table 10. From these results, we can observe that the disparity of higher-order critiquing ability exists in specific models. Although performance is limited, Qwen-2.5-14B-Instruct exhibits more effective in high-order critiquing, as we can observe progressive improvements between response and recursive critiques across stages. Further performance gap analysis in Table 11 reveals that the main discrepancies may arise from those gap in distinguishing the true statements from mixed true and false inputs.
@@
Prompt for $C^{2}$ Generation You are given a multiple-choice question. And two individuals, Person 1 and Person 2, have selected their answers and provided their explanations for their choices. Additionally, two more individuals, Reviewer 1 and Reviewer 2, have read these explanations and provided their evaluations of Person 1’s and Person 2’s reasoning. Your task is to identify which answer is correct based on their explanations and the information known about the question. Follow this structure for your response: - Explanation of Choice: Compare both explanations to your knowledge about the topic and determine which aligns better with the correct answer, starts with ‘Explanation: ‘. - Analysis of Other Options: Review the explanation provided by each person. Evaluate the reasoning and evidence behind each choice and point out any inaccuracies or correct assumptions. - Answer: On a separate line, starts with ‘Answer: ‘, state your chosen option (A, B, C, or D) only, without any additional text. ### Question: {question} ### Options: {options} ### Person 1’s Response: {gen1} ### Person 2’s Response: {gen2} ### Reviewer 1’s Response: {c1} ### Reviewer 2’s Response: {c2}
Figure 12: AI generartion template in $C^{2}$ Stage (Part 1)
Prompt for $C^{2}$ Generation Example Input: ### Question: Which element has the atomic number 6? ### Options: A) Nitrogen  B) Oxygen  C) Carbon  D) Helium ### Person 1’s Response: Chosen Answer: C Explanation: Carbon is the element with atomic number 6, well-known for being the basis of organic chemistry. ### Person 2’s Response: Chosen Answer: A Explanation: Nitrogen is important for life on Earth, making up a large portion of the atmosphere. ### Reviewer 1’s Response: Explanation: Person 1 accurately states that Carbon has the atomic number 6. This explanation is correct and directly addresses the question. Carbon is indeed a fundamental element … Answer: C ### Reviewer 2’s Response: Explanation: Person 1 has correctly identified Carbon as the element with atomic number 6. Their … Answer: C Example Output: Explanation: - Reviewer 1 provides a clear and direct evaluation of Person 1’s and Person 2’s responses. They … Answer: C
Figure 13: AI generartion template in $C^{2}$ Stage (Part 2)
@@
Appendix D Broader Impacts
Our recursive self-critiquing framework offers significant potential for maintaining effective AI oversight as capabilities surpass human abilities, potentially democratizing participation in AI governance through more tractable evaluation processes. However, this approach also carries risks, including the possibility of creating false confidence in oversight effectiveness given current limitations in AI critique capabilities, vulnerability to adversarial examples designed to bypass critique chains, and concerns about computational resource requirements potentially limiting access to effective oversight tools among less-resourced stakeholders. We acknowledge these potential impacts and encourage continued research to strengthen the reliability of scalable oversight methods.